LBC Tech Chip IA Commissioning - 08 May 2010 UT

Night 4

Observers: M.Pedani (LBTO)
Telescope Operator: SAllanson (LBTO), JHill (LBTO)
Telescope Support: JUrban(LBTO)
Instrument Support: DOfficer (LBTO)
LBC team: RSpeziali (INAF), ABaruffolo(INAF)

Summary

Nautical Twilight at 03:04 UT

Tonight we could do more tests of the Tech Chip IA algorithm while taking some good quality science data. The thick cirrus clouds arrived around 8:00 UT and we had to switch to Andrew tests (collimation + pointing).

A.Baruffolo did some last changes this afternoon before starting observations. With respect to yesterday, he changed the code to take into account the fact that, at Blue, the pupils are slightly cut on one side and not perfectly circular.

I haven't analyzed the images yet because we are almost closing the dome but, after looking almost all the night at the corrections that were applied, we are confident that the algorithm is working fine. We could see some big Z4 correction but only in correspondence of significant changes in Z11. Other than that, the corrections sent are no more than 400-600nm.

So, the plan for tomorrow (first LBTB night) is to leave the code in place. Obviously the best results will be obtained if a suitable star is always placed in the Tech Chip.

Also, I put on LBC the new Filters Focus offsets on both LBCB and LBCR that I determined the first night.

The first set of SX pointing data with Andrew's collimation model was extrafocal. We fixed that, but only got part of a second set.

Preparations

Updated DXPMLBCCollimation.dat to remove the DX coma that we found last night.
The new DX M1 Global Offsets are: X=-0.065 Y=-0.633 Z=1.422 RX=-15.742 RY=-10.487
Note that this removes the big Xglobal that I had to add on 17Feb. It remains a mystery as to what caused that large offset in the first place.

Changed Filters Focus Offsets on lbccontrol with values obtained on First Tech Night.; Old config files appended " _20100508"

Open

Warmer than yesterday

02:09 Photometric, T_out = 8.3C Humidity =28 % T_mirrorSX=6.1C, T_mirorr_DX=6.0C; Wind up to 17m/s

Details

02:46 still fighting with Firefox on obs2 machine; S.Hooper on the phone

03:03 LBCR RPC Client/Host Error

03:04 problem still there. I open Firefox from lbcobs account on obs2

03:12 Slewing to WT10_292 to check Pointing; CA=-76.8 IE=-19.7

03:19 Slewing to WT10_086 to Low Elevation to check IA; IA=-384.2

03:28 Slew to Sci Field Cosmos Fiore_UspecZ_P3_c and DOFPIA; Blue starts with Z4=-8887nm

03:57 OB stopped; Blue shows pupils; we find an error in the LBCB Filters offsets configuration file
May  8 03:57:29 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -209 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:00:31 tcs3 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:01:05 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 111 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:04:03 tcs3 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:04:05 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 326 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0

04:10 Same Field; recollimate
May  8 04:20:30 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:24:03 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 208 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:24:16 tcs3 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:27:50 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -339 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:30:54 tcs3 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:31:26 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -287 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:35:02 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 189 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:38:38 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -183 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:42:16 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 100 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:45:20 tcs3 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 -445 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:45:52 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -50 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:49:28 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -267 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 04:53:05 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 43 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0

04:56 Slew to Sci Field Cosmos Fiore_UspecZ_P3_d; Z4 corrections are applied correctly
May  8 04:58:51 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 05:01:59 tcs3 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 05:02:26 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 471 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 05:06:04 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -502 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 05:09:08 tcs3 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 05:09:43 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -214 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 05:16:25 tcs3 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 05:20:27 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -528 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 05:24:04 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -209 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 05:27:41 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -80 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 05:30:47 tcs3 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 -775 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 05:31:18 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -38 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
05:23 T_out = 7.6, RH = 29%, T_mirrorSX=7.9 C, T_mirorr_DX=7.7 C; Wind up to 10m/s

05:33 Slew to Sci Field N5053_f1_1 and DOFPIA
May  8 05:49:20 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 05:49:34 tcs3 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 05:56:44 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 103 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 05:57:06 tcs3 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 129 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 06:01:56 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 06:05:31 tcs3 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 06:05:49 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -110 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 06:09:45 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -124 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 06:13:03 tcs3 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 139 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
May  8 06:13:39 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -67 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0

05:58 Slew to Sci Field N5053_f1_2

06:15 DOFPIA

06:22 Slew to Sci Field N5053_f1_3

06:39 Slew to Sci Field N5053_f1_4

06:56 Slew to Sci Field N5053_f1_5

07:13 DOFPIA

07:19 Slew to Sci Field N5053_f2_1

07:54 Slew to Sci Field N5053_short_1

08:03 DOFPIA

08:06 Slew to Sci Field N5053_short_2

08:14 Slew to Sci Field N5053_short_3

08:22 Slew to Sci Field N5053_field

08:35 STOPPED due to CLOUDS

Pointing Data for Andrew

08:40 We decide to do Andrew stuff with a new pointing model and Coll. Table

Andrew's Procedure:
Subject: Pointing data procedure
Date: Thu, 6 May 2010 13:00:10 -0700

This is an outline of the procedure I'd like you to use to get pointing data
for me. All of this is for LBC blue, I'm not interested in any red data at
this stage.

0) P.S. Turn on pointing logging

1)      Using the usual LBC collimation model, DOFPIA on a high elevation
star once converged note the total collimation value for Z on SX M1.


Missing step: Remove the -0.8 mm instrument offset before recording Ztotal.


2)      Switch collimation models to the new SX M1 collimation model that
John has prepared for this task.


Missing step: Clear the active optics collimation corrections.


3)      Adjust the SX M1 PSF Global Offset value for Z to make the "Total
Collimation" value for Z identical to that recorded in step 1.

4)      Take pointing data for 28 stars.

5)      Check the images, there will be bad (very bad) coma and maybe
exposure times will need to be adjusted as a result, best to check them as
you go along I guess.

08:41 slewing to ACT0292 EL=+82 and DOFPIA

08:45 Started Log for Pointing; filename = 20100508UT.ptlog

08:52 SX Ztotal=-0.942 mm. We load new Coll Table (from Andrew's laser tracker data) and we get Ztotal=-0.594; Global offset applied of Zglobal=-0.370 (to get us back to -09.42) and we start taking Pointing Data

Images are out of focus (extrafocal pupils) because the -0.8mm offset remained after DOFPIA was finished.
Also we didn't clear the active optics corrections from DOFPIA.

08:59 Slewing to ACT0369
09:08 Slewing to WT10_400
09:11 Slewing to WT10_399
09:16 Slewing to WT10_354
09:20 Slewing to ACT_292
09:25 Slewing to WT10_173
09:28 Slewing to WT10_247
09:31 Slewing to WT10_092
09:34 Slewing to ACT_162
09:36 Slewing to ACT_268
09:39 Slewing to ACT_202
09:42 Slewing to WT10_321
09:45 Slewing to ACT_334
09:47 Slewing to ACT_289
09:50 Slewing to WT10_255
09:52 Slewing to ACT_179
09:55 Slewing to WT10_094
09:58 Slewing to WT10_109
10:00 Slewing to ACT_300
10:03 Slewing to WT10_264
10:06 Slewing to WT10_207
10:09 Slewing to WT10_311
10:12 Slewing to ACT_330
10:15 Slewing to WT10_396
10:17 Slewing to ACT_404
10:20 Slewing to ACT_411
10:22 Slewing to ACT_345
10:25 Slewing to ACT_302

End 28 stars (pupils)!

10:33 Ztotal = -0.814 at EL=79.0

Adjust SX Zglobal -0.578 and clear active optics corrections.
Add +0.8 to Zglobal making it 0.222 mm.

10:37 repeat ACT_302 (now with in-focus stars that are only slightly comatic)

Testing gopoint_lbc.cl and collecting some in-focus data for Andrew

Note the attached log file below, but also note that it has discrepancies.

10:40 gopoint_lbc 29 ACT0382

10:42:42 Another LBC-Red RPC error

Turn Off Cameras and lbckill/lbcstart; Turn on (lost 10 minutes)

10:52 repeat the gopoint_lbc OB on ACT0382 manually. Note that the OBJECT NAME is not in the header (worse: the wrong name is in the header) because LBC got a generic OB from gopoint_lbc.cl.

10:55 gopoint_lbc 30 ploop+ first object is WT10_375

10:57 Bug in my script didn't know how to wait for the OB to finish. So WT10_375 is trailed by the next preset.

11:01 gopoint_lbc 30 ploop+ - Now the script sleeps 60 sec after playing OB. WT10_386

11:05 WT10_405
11:07 WT10_406
11:09 WT10_369

11:12 gopoint_lbc 34 ploop+ WT10_385 (two presets collided so it is trailed - not sure what made the collision, but suspect it was a 4AM problem with the driver) Sleep was increased to 80 sec, but still there are issues.

11:15 ACT0354
11:17 WT10_367
11:19 ACT0315 (trailed?) but then another image at 11:21 ???

11:23 gopoint_lbc 38 ploop+ ACT0300 (unwrapping azimuth - 300 sec timeout on the preset)

11:29 WT10_211
11:33 WT10_175
11:36 WT10_160
11:38 WT10_179

Note that these images do NOT have horrible coma, so this laser tracker data is a good start of an LBC-Blue collimation model.

Ran out of darkness before we could finish a full set of in-focus data.

Close

11:41 Close for impending dawn T=6 D=-8 Wind 5m/sec. Sky partly cloudy.

11:44 Put back the normal LBC-Blue collimation model. (11-Sep-2009)

-- JohnHill - 07 May 2010
I Attachment Action Size Date Who Comment
gopoint20100508.loglog gopoint20100508.log manage 1 K 10 May 2010 - 22:00 JohnHill Editted log file from the gopoint_lbc tests
gopoint_lbc.clcl gopoint_lbc.cl manage 4 K 10 May 2010 - 21:15 JohnHill Script for automatic LBC pointing data (prototype)
Topic revision: r8 - 10 May 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback